Solve the system by the addition method:
Eq 1 states: 3x = 3 + 2y
Eq 2 states: -4/3x + y = 1/3

Answer by stanbon(75887)

Solve the system by the addition method:
Eq 1 states: 3x = 3 + 2y
Eq 2 states: (-4/3)x + y = 1/3
------------------
Rearrange Eq 1 and multiply thru Eq 2 by 3 to get:
#3: 3x-2y=3
#4: -4x+3y=1
--------------
Multiply #3 by 3 and #4 by 2 to get:
#5: 9x-6y=9
#6:-8x+6y=2
-------------
Add to get:
x = 11
---------
Substitute into #3 to solve for y:
3*11-2y=3
2y = 30
y = 15
------------
Check the answer in -4x+3y=1
-4*11 + 3*15 = 1
-44+45 = 1
1=1
============
Cheers,
Stan H.